OpenStudy (zyberg):

Is there any simple explanation of why LC circuit period is equal to square root of LC?

OpenStudy (zyberg):

@ganeshie8 maybe you would have any ideas?

OpenStudy (osprey):

I think that it's to do with the way the differential equation or the LC - or any other oscillatory circuit - is set up. in SHM, it's x double dot = - omega SQUARED x This leads to square roots in just about everything oscillatory - re the formula for a simple pendulum and it's periodic time, eg. http://perendis.webs.com

OpenStudy (irishboy123):

that's the angular freq, not the period.....but ... really depends on what you mean by "why" you can just solve the DE: \(\ddot I +\omega _{0}^{2}I =0\), and you're coming at it like you know all that. ....and so, then you can say that "that's that". but what is "why". https://www.youtube.com/watch?v=36GT2zI8lVA i can't see a physical analogue, even from a simple spring or pendulum system. i'm guessing that's what you're asking cos your question is so general. sorry can't be more helpful.

OpenStudy (osprey):

come to think of it, it's probably another "mathematical device", shall I say, to make things tidier/easier further down the chain. MAY be a bit like doing ict as the "fourth dimension" ... ct "converts" speed and time into "space-time" - distance, and the i (root of -1 ie imaginary) makes things look tidy in some invariant or other in spec rel. And, after all, what is "i" but the square root of -1. Imaginary. Some people's imaginations ...

OpenStudy (osprey):

@IrishBoy123 I thought I saw a post about metre in poetry a day or so ago. It seems to have vanished both open and closed can't find it anywhere. Is this my imagination, or did I really see Iambic Pentameter ?????

OpenStudy (radar):

you could look at it like this: Period (time) = 1/f frequency of an LC circuit (resonance):\[f=1/(2\pi \sqrt{LC})\]You can now write it for Period as:\[2\pi \sqrt{LC}\] Why you may ask? Check out the link.

OpenStudy (osprey):

@radar Why you may ask? .... looks like the answer is to make the maths elegant. After all, doesn't maths/physics delight in straight line graphs and any graph that can be "curve fitted" ? So, the practical people do the "knob twiddling", hand the results to the analysts who do the graphs and get the knob twiddles to match the maths and get the publications ?

OpenStudy (irishboy123):

Trying out an anology ..... why is the period of a mass-spring system given by \(2 \pi \sqrt {\dfrac{m}{k}} \) ?? We know from energy conservation that: \(\frac{1}{2} k x^2 + \frac{1}{2} m \dot x ^2 = const \qquad \circ\). Why is energy conserved? Well, for present purposes, it just is. Why is \(U_s = \frac{1}{2} k x^2\)? We have Hooke's Law: \(\frac{dF}{dx} = k\), and we suppose that the intermolecular forces are inverse square but they linearise over short distances giving this linear relationship. We then have all of Newton's ideas/definitions, especially that \(W \equiv E \equiv F \cdot x\), and so we get that \(\frac{1}{2} k x^2\) idea pretty easily. Same for \(T = \frac{1}{2}m\dot x ^2\) But if we then plot the phase portrait \(x\) vs \(\dot x\) we have an ellipse ...... and that seems to be the key observation Specifically, from \(\circ\): \(\left( \dfrac{x}{\sqrt{m}} \right)^2 + \left( \dfrac{\dot x}{\sqrt{k}} \right)^2 = const \), clearly an ellipse Or \(\left( \dfrac{x}{\gamma \sqrt{m}} \right)^2 + \left( \dfrac{\dot x}{\gamma \sqrt{k}} \right)^2 = 1 \qquad \triangle\) .....where \(\gamma\) is the interim fudge factor... We can parameterise any ellipse as \(x = a \cos \omega t, y = b \sin \omega t \), ie: \(\left( \dfrac{x}{a} \right)^2 + \left( \dfrac{y}{b} \right)^2 = 1 \qquad \star\) and we also know from calculus that in our phase portrait with our particular ellipse that \(y = \dot x\) and \(y = \dot x = - a \omega \sin \omega t\), where \(\omega\) can be any angular frequency But we can pattern match \(\star\) to: \(\left( \dfrac{a \cos \omega t}{a} \right)^2 + \left( \dfrac{- a \omega \sin \omega t}{b} \right)^2 = 1 \qquad \) ie \(\omega = \dfrac{b}{a}\) Then from \(\triangle\) we have only one conclusion: \(\omega = \dfrac{\gamma \sqrt{k}}{\gamma \sqrt{m}} = \sqrt{ \dfrac{k}{m} }\) So the answer in the case of a spring-mass system seems to be that the law of conservation of energy, Hooke's and Newton's Laws, calculus and the geometry of an ellipse collectively demand that \( T = 2 \pi \sqrt{ \dfrac{m}{k} }\) So, after all that faffing about, the smart answer I think is that the DE says it has to be so. As it does, mutatis mutandis, in the case of LC.

OpenStudy (radar):

@osprey Yes, not only in order to make the math elegant, but the speaker also. lol

OpenStudy (osprey):

@IrishBoy123 mutatis mutandis ... that looks suspiciously like Latin ? .....................................

OpenStudy (irishboy123):

@osprey good spot !!!! you do the French, i'll do the Latin ?!?!

OpenStudy (osprey):

@IrishBoy123 What does that latin mean ?

OpenStudy (irishboy123):

literally, and Latin is beautifully concise, it means "all that needs to have been changed having been changed". so in an equation, you will have moved subscripts around. in plain English, i think the nearest idiom is "all things being equal". but i don't really get that expression.